Circular No. 6768

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dp IN NGC 969 L. Arnold, Observatoire de Haute-Provence, reports the discovery of

an apparent supernova by Denis Christen, Hesingue, France, on unfiltered CCD images taken with a 0.3-m telescope on Nov. 1.10 and 1.93 UT, at magnitudes 17.8 ± 0.4 and 17.9 ± 0.4, respectively. The supernova is located at α = 2^{h}34^{m}06^{s}.48, δ = +32^{o}56′13′′.4 (equinox 2000.0), which is 22′′ west and 36′′ south of the center of NGC 969. The object is absent from several photographic atlses, including the Digital Sky Survey. The Nov. 1.10 image (the sum of four 3-min frames; limiting mag ∼ 20) is available at

http://www.obs-hp.fr/~arnold/NGC969.JPG. G. M. Hurst, Basingstoke,

England, reports that prediscovery images of SN 1997dp (measured position end figures 06^{s}.48, 13′′.8) were recognized by M. Armstrong, Rolvenden, Kent, on a ‘patrol master’ frame obtained on Sept. 10.1 (mag 16), as well as on frames taken on Sept. 28 and Oct. 14 (mag ∼ 17).

S. Jha, P. Challis, P. Garnavich, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that SN 1997dp is a type-Ia supernova, ∼ 6 weeks past maximum, based on a spectrum taken by Jha and Challis at the Multiple Mirror Telescope on Nov. 4.2 UT. The spectral features of the supernova are consistent with the de Vaucouleurs’ catalogue recession velocity for NGC 969 of 4500 km/s.

η CARINAE

Z. Abraham, Instituto Astronômico e Geofísico (IAG), Universidade de São Paulo (USP); and A. Damineli, IAG/USP and Joint Institute for Laboratory Astrophysics, Boulder, report the fading of the 7-mm flux density of η Car by a factor of three since its maximum in 1995: “Our observations were made with the Itapetinga radiotelescope in Brazil. The measured flux density was 3.6 ± 0.3 Jy. The He I 1083.0-nm line is fading as predicted (Damineli 1996, Ap.J. 460, L49) by the 5.52-yr period. The high-excitation lines are expected to fade completely around Dec. 10 through Jan. 20.”

COMET 43P/WOLF-HARRINGTON Total visual magnitude estimates by R. J. Bouma, Groningen, The

Netherlands (0.25-m reflector): Oct. 11.14 UT, 12.7; 28.11, 12.8; 31.16, 12.6; Nov. 4.17, 12.8.

1997 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green